The inverse demand for a product is P=50-0.1Q, and the mc to produce is $10. What is the best pricing strategy for this firm: charging all customers the same price, 2-part pricing, or block pricing?
Answer #1

The form has to charge same price to all customers because only one demand curve exists for this product.

P=50-0.1Q

TR=PQ

TR=50Q-0.1Q^2

MR=d(TR)/dQ

MR=50-0.2Q

MC=10

At equilibrium MR=MC

50-.2Q=10

.2Q=40

Q=200

P=50-.1(200)

P=30

Firm charges $30 to all customers